# Mode analysis of Nambu-Goldstone modes in U(1) charged first-order relativistic viscous hydrodynamics

### v2

@@ 3.1 Symmetry structure in hydrodynamics
Let us consider the following symmetry transformations for the external fields in the path integral Eq. (8):
  1. 1.
    the physical spacetime diffeomorphism, i.e., general coordinate transformations generated by a tangent $\xi^{\mu}(x)$:
    $x^{\mu}\to x^{\mu}+\xi^{\mu}(x).$ (136)
  2. 2.
    the noise diffeomorphism generated by a noise tangent $\bar{\xi}^{\mu}(x)$:
    $X$ $\to X+\mathcal{O}(\hbar^{2}),$ (137)
    $\bar{X}$ $\to\bar{X}+\pounds_{\bar{\xi}}X+\mathcal{O}(\hbar^{3}),$ (138)
    where $X=g_{\mu\nu},A_{\mu},\cdots$.
  3. 3.
    the gauge transformation:
    $A_{\mu}\to A_{\mu}-\nabla_{\mu}\Lambda,$ (139)
  4. 4.
    the noise gauge transformation:
    $\bar{A}_{\mu}\to\bar{A}_{\mu}-\nabla_{\mu}\bar{\Lambda}.$ (140)
With this symmetry structure, the generating functional (8) can be regarded as a local effective action of hydrodynamics. In fact, the invariance under the noise gauge transformation and the noise diffeomorphism yields the equation of motions. This fact implies that the parameters of these transformations are regarded as the (noise) dynamical variables. Using the Stueckelberg trick, one may introduce the NG modes by
$\xi^{\mu}\to-\pi^{\mu},~{}\bar{\xi}^{\mu}\to-{\bar{\pi}}^{\mu},~{}\Lambda\to-\theta~{}\bar{\Lambda}\to-\bar{\theta}.$ (141)
Then, as expected, the equations of motion are found as
$\frac{\delta I}{\delta\bar{\theta}}=0$ $\to\nabla_{\mu}J^{\mu}=0,$ (142)
$\frac{\delta I}{\delta{\bar{\pi}}_{\mu}}=0$ $\to\nabla_{\mu}T^{\mu\nu}+J_{\mu}F^{\mu\nu}=0.$ (143)
Next, let us write $T$, $J$, $\cdots$ by the physical NG modes. By construction of Eq. (141), the NG modes transform as follows:
$\pi^{\mu}(x)$ $\to\pi^{\prime\mu}(x^{\prime})=\pi^{\mu}(x)+\xi^{\mu}(x),$ (144)
$\theta(x)$ $\to\theta^{\prime}(x)=\theta(x)+\Lambda(x).$ (145)
With $\pi^{\mu}$ and $\theta$, one finds a set of four scalar fields and an invariant phase:
$\sigma^{\alpha}=x^{\alpha}-\pi^{\alpha},\phi=\varphi-\theta,$ (146)
where $\varphi$ is the global phase. $(\sigma^{\alpha},\phi)$ are regarded as an internal spacetime of fluid labels (fluid spacetime) with the following internal symmetries that “define” a fluid Crossley:2015evo :
  1. 1.
    Time independent spacetime diffeomorphism:
    $\sigma^{\alpha}\to\sigma^{\alpha}+f^{\alpha}(\sigma^{b}),$ (147)
    which implies that one can define a spatial fluid label once, and a constant time hypersurface is fixed. One may allow spacetime-dependent time diffeomorphism in compensation for the ambiguity in the constant time hypersurface and, thus, the temperature. In this work, we fix this freedom implicitly.
  2. 2.
    Spatial dependent phase shift:
    $\phi\to\phi+f_{4}(\sigma^{a}),$ (148)
    which implies that we can fix the phase of each fluid element once.
The hydrodynamical theory is constructed from fluid spacetime invariants. Firstly, one may introduce an invariant 4-vector $\beta^{\alpha}=\beta_{\rm c}\delta^{\alpha}_{0}$ in fluid spacetime, with a scaling factor $\beta_{\rm c}$. Note that the superscript $\alpha$ implies a tensor in fluid spacetime, which is distinct from physical spacetime index $\mu$. One can readily check that $\beta^{\alpha}$ is invariant under Eq. (147). One may embed $g_{\mu\nu}$ and $A_{\mu}$ into the fluid spacetime by regarding Eq. (146) as an effective coordinate/phase transformation. We find
$h_{\alpha\beta}$ $=\frac{\partial x^{\mu}}{\partial\sigma^{\alpha}}\frac{\partial x^{\nu}}{\partial\sigma^{\beta}}g_{\mu\nu},$ (149)
$B_{\alpha}$ $=\frac{\partial x^{\mu}}{\partial\sigma^{\alpha}}A_{\mu}+\frac{\partial}{\partial\sigma^{\alpha}}(\theta+\phi).$ (150)
Using $h_{\alpha\beta}$, we can introduce the following scalar field in fluid spacetime:
$\beta=\sqrt{-\beta^{\alpha}\beta_{\alpha}}=b\beta_{\rm c},~{}b\equiv\sqrt{-h_{00}},$ (151)
which is identified with the inverse temperature of the fluid element $\sigma^{\alpha}$. The 4-velocity of the fluid is identified with
$u^{\alpha}=\frac{\beta^{\alpha}}{\beta}=\frac{\delta^{\alpha}_{0}}{b}.$ (152)
$B_{\alpha}$ is not invariant under Eq. (148). One can construct the following gauge-invariant scalar and tensor:
$\mu=u^{\alpha}B_{\alpha},~{}F_{\alpha\beta}=\nabla_{\alpha}B_{\beta}-\nabla_{\beta}B_{\alpha},$ (153)
where $\mu$ is identified with the chemical potential of the fluid element $\sigma^{\alpha}$. One may consider the pull-backs of these hydrodynamical variables in the physical spacetime by using Eq. (146), and then $\pi^{\mu}$ and $\theta$ are embedded into physical spacetime. These NG modes are the “slow modes” that are distinct from the fast mode integrated in Eq. (8). The slow modes are long-lived collective excitations around hydrodynamical solutions, which were absent when writing Eq. (8). One may also integrate the NG modes. The integration over the additional long-lived excitation will introduce nonlocal terms in Eq. (8).
Similarly, the covariant noise fields are constructed as follows. The noise diffeomorphism and gauge transformations are
$\delta_{\bar{\xi},\bar{\Lambda}}g_{\mu\nu}$ $=0,$ (154)
$\delta_{\bar{\xi},\bar{\Lambda}}\bar{g}_{\mu\nu}$ $=\nabla_{\mu}\bar{\xi}_{\nu}+\nabla_{\nu}\bar{\xi}_{\mu},$ (155)
$\delta_{\bar{\xi},\bar{\Lambda}}A_{\mu}$ $=0,$ (156)
$\delta_{\bar{\xi},\bar{\Lambda}}\bar{A}_{\mu}$ $=-\nabla_{\mu}\bar{\Lambda}+\bar{\xi}^{\nu}\nabla_{\nu}A_{\mu}+A_{\nu}\nabla_{\mu}\bar{\xi}^{\nu}.$ (157)
Then, one may introduce the following noise diffeomorphism and gauge invariants:
$\bar{G}_{\mu\nu}$ $=\bar{g}_{\mu\nu}-\nabla_{\mu}{\bar{\pi}}_{\nu}-\nabla_{\nu}{\bar{\pi}}_{\mu},$ (158)
$\bar{C}_{\mu}$ $=\bar{A}_{\mu}+\nabla_{\mu}\bar{\theta}-{\bar{\pi}}^{\nu}\nabla_{\nu}A_{\mu}-A_{\nu}\nabla_{\mu}{\bar{\pi}}^{\nu}.$ (159)
To summarize, $\Phi=\{\beta,u^{\mu},\mu,F_{\mu\nu},g_{\mu\nu}\}$, $\bar{\Phi}=\{\bar{G}_{\mu\nu},\bar{C}_{\mu}\}$ and their covariant derivatives are the symmetric building blocks of the dynamical action of the NG modes.
@@
When NG modes are set to zero, the effective action must reproduce the UV generating functional (8). Hence, the effective action for the NG modes should be written as Glorioso:2017fpd
$I[\Phi,\bar{\Phi}]=\int d^{4}x\sqrt{-\det g}\sum_{n=1}\mathcal{L}_{n},$ (160)
@@
$\mathcal{L}_{1}$ $=\frac{1}{2}T^{\mu\nu}[\Phi]\bar{G}_{\mu\nu}+J^{\mu}[\Phi]\bar{C}_{\mu},$ (161)
$\mathcal{L}_{2}$ $=\frac{i}{4}W^{\mu\nu\rho\sigma}[\Phi]\bar{G}_{\mu\nu}\bar{G}_{\rho\sigma}+iX^{\mu\nu\rho}[\Phi]\bar{G}_{\mu\nu}\bar{C}_{\rho}+iY^{\mu\rho}[\Phi]\bar{C}_{\mu}\bar{C}_{\rho}.$ (162)
@@
3.2 Effective action for Nambu-Goldstone modes
For simplicity, let us consider the following simplest background fields:
$g_{\mu\nu}=\eta_{\mu\nu},~{}A_{\mu}=\mu_{\rm c}\delta^{0}_{\mu}.$ (163)
That is, we consider a Minkowski background for physical spacetime, and we only have a global phase and zero field strength. The Minkowski metric is defined as $\eta_{\mu\nu}={\rm diag}(-1,1,1,1).$ From Eq. (151), one finds
$b=\sqrt{-h_{00}}=\sqrt{-\frac{\partial x^{\mu}}{\partial\sigma^{0}}\frac{\partial x^{\nu}}{\partial\sigma^{0}}g_{\mu\nu}}=\sqrt{(1+\dot{\pi}^{0})^{2}-\dot{\pi}^{i}\dot{\pi}^{j}\delta_{ij}}=1+\dot{\pi}^{0}+\mathcal{O}(\partial\pi^{2}).$ (164)
Hereafter, we truncate the nonlinear term in NG modes. Then, the temperature field in the physical spacetime is written as
$\beta=\beta_{\rm c}(1+\dot{\pi}^{0}).$ (165)
The spatial NG mode now perturbs the fluid velocity:
$u^{\mu}(x)=\frac{\partial x^{\mu}}{\partial\sigma^{\rho}}b^{-1}\delta^{\rho}_{0}=(1-\dot{\pi}^{0})\delta^{\mu}_{0}+\dot{\pi}^{\mu}=(1,\dot{\pi}^{i}).$ (166)
Similarly, the chemical potential is found as
$\mu=u^{\mu}A_{\mu}+u^{\mu}\nabla_{\mu}(\theta+\phi)=\mu_{\rm c}+\dot{\theta}.$ (167)
Thus, while $\mu$, $\beta$, and $u^{\mu}$ are zeroth order in the constitutive relations, they are first order with respect to the derivatives on the NG modes. The induced metric $\gamma^{\mu\nu}$ is defined with respect to $u^{\mu}$, which is distinct from the induced metric with respect to the constant time hypersurface. We find
$\gamma^{00}=0,~{}\gamma^{0i}=\dot{\pi}^{i},~{}\gamma^{ij}=\delta^{ij}.$ (168)
Hydrodynamical coefficients are perturbed via $\beta$ and $\mu$. For example, $\rho$ is expanded into
$\rho(\beta,\mu)=\rho_{\rm c}+\beta_{\rm c}\rho_{\beta}\dot{\pi}^{0}+\rho_{\mu}\dot{\theta},$ (169)
where $X_{\mu}=(\partial X/\partial\mu)_{\beta}$ and $X_{\beta}=(\partial X/\partial\beta)_{\mu}$ for $X=P,\rho$. Hereafter, the hydrodynamical coefficients are always expanded in the NG modes, so we omit “c” for notational simplicity. That is, we redefine the symbols as $\beta\equiv\beta_{\rm c}$ and $\mu\equiv\mu_{\rm c}$.
@@ 3.2 Effective action for Nambu-Goldstone modes
$I_{s}^{(1,0,0)}=\frac{1}{2}\int d^{4}x\Bigg{[}\rho\bar{g}_{00}+P\delta^{ij}\bar{g}_{ij}+2n\bar{A}_{0}\Bigg{]}.$ (171)
@@
$\bar{A}_{0}(n_{\mu}\dot{\theta}+\beta n_{\beta}\dot{\pi}^{0})+\bar{A}_{i}(n\dot{\pi}^{i})\Bigg{]}.$ (172)
@@
$+\bar{A}_{i}\beta\left(-\beta\chi_{1}\left(\partial^{i}\dot{\theta}+\mu\partial^{i}\dot{\pi}^{0}\right)+\tau_{1}\partial^{i}\dot{\pi}^{0}-\tau_{2}\ddot{\pi}^{i}\right)\Bigg{]}.$ (173)
By differentiating Eq. (173), one finds the first order fluctuations to $T$ and $J$. The noise term is already presented in Eq. (70).
@@
$I_{d}^{(1,0,0)}=-\int d^{4}x\Bigg{[}(\rho-\mu n)\dot{{\bar{\pi}}}_{0}+P\partial^{i}{\bar{\pi}}_{i}-n\dot{\bar{\theta}}\Bigg{]}.$ (174)
@@
$I^{(1,0,0)}_{d}=\int d^{4}x\Bigg{[}(\dot{\rho}-\mu\dot{n}){\bar{\pi}}_{0}+\dot{n}\bar{\theta}\Bigg{]}.$ (175)
@@
$+n\partial_{i}\bar{\theta}\dot{\pi}^{i}+\dot{\bar{\theta}}\left(\dot{\theta}n_{\mu}+\beta\dot{\pi}^{0}n_{\beta}\right)\Bigg{]}.$ (177)
@@
$+\partial_{i}\bar{\theta}\beta\left(-\beta\chi_{1}\left(\partial^{i}\dot{\theta}+\mu\partial^{i}\dot{\pi}^{0}\right)+\tau_{1}\partial^{i}\dot{\pi}^{0}-\tau_{2}\ddot{\pi}^{i}\right).$ (178)
@@
$-2\nu_{1}\dot{\bar{\pi}}_{0}(\dot{\bar{\theta}}+\mu\dot{\bar{\pi}}_{0})-2\tau\left(\partial_{i}{\bar{\pi}}_{0}+\dot{{\bar{\pi}}}_{i}\right)(\partial^{i}\bar{\theta}+\mu\partial^{i}\bar{\pi}_{0})-2\nu_{2}(\dot{\bar{\theta}}+\mu\dot{\bar{\pi}}_{0})\partial^{i}\bar{\pi}_{i}\Bigg{]}.$ (179)

### v3

@@ 4.1 Symmetry structure in hydrodynamics
The structure of SSB in hydrodynamics is more intricate than in the vacuum case due to the doubling principle inherent in the Schwinger–Keldysh formalism. Recent developments in the EFT perspective on hydrodynamics suggest that hydrodynamical modes are NG modes associated with so-called strong-to-weak spontaneous symmetry breaking (SWSSB) [18]. Accordingly, we refer to $\pi^{\mu}$ as the NG mode in the context of SWSSB throughout this work.
The introduction of an NG mode for SSB of a global $U(1)$ scalar field theory proceeds as follows. We parameterize the complex scalar field as $\Phi=re^{i\varphi}$. The global $U(1)$ symmetry is spontaneously broken by the vacuum. Excitations around a chosen vacuum phase $\varphi=\varphi_{0}$ are described by a field $\theta$, which nonlinearly realizes the symmetry broken by the vacuum via a constant shift $\theta\to\theta+\epsilon$. The radial mode $r$ is typically a heavy degree of freedom that is integrated out in hydrodynamics. Introducing $\theta$ leads to a shift in the external background field: $A_{\mu}\to A_{\mu}+\nabla_{\mu}\theta$. This transformation is distinct from a gauge transformation, since the $U(1)$ symmetry is global and $A_{\mu}$ is a nondynamical background field.
This structure is not specific to internal symmetry breaking. Consider a theory of a real scalar field $\phi$ in Minkowski spacetime. Suppose the dynamics admits a time-dependent solution $\phi=\bar{\phi}(t)$. Although the action is invariant under time translations, the solution breaks this symmetry spontaneously. The excitation around the background solution can then be described by an NG mode $\pi(x)$ via
$\phi(x)=\bar{\phi}(t+\pi(x)).$ (4.1)
Under time translations, the field $\pi(x)$ transforms with a constant shift, reflecting the nonlinearly realized time translation symmetry. These NG modes can be understood as reparametrizations of the original field variables, defined with respect to specific background solutions. More precisely, they correspond to the exponential map generated by the broken symmetry generators acting on the condensates that spontaneously break the symmetry.
The same idea can be applied to hydrodynamics. The local thermal vector $\beta^{\mu}$ is now reparametrized as
$\beta^{\mu}=e^{-\pounds_{\pi}}\beta^{\mu}_{c},$ (4.2)
where $\beta^{\mu}_{c}=\beta_{c}\delta^{\mu}_{0}$ with a constant $\beta_{c}$, and £$\pounds_{\pi}$ denotes the Lie derivative along $\pi^{\mu}$. Eq. (4.2) represents a local thermal state as a diffeomorphic embedding of a global thermal state.
For global symmetries, the NG modes have a clear physical meaning as excitations around the broken vacuum. When gauging the symmetries, one can embed $\pi^{\mu}$ in the same way as in the global symmetry case; however, the physical meaning of the hydrodynamical modes depends on the choice of gauge. In the conventional formulation [15], the NG mode is eaten by the metric in the unitary gauge, where the fluid is unperturbed and the hydrodynamical theory is constructed. The fluid is then defined by additional symmetry principles, with the thermal vector $\beta^{\mu}$ introduced as an emergent background quantity.
In contrast, in our approach we begin by specifying the physical thermal state through a given vector $\beta^{\mu}$, which encodes the local rest frame and temperature distribution. We introduce $\beta^{\mu}$ in a gauge where the metric perturbations are eliminated, corresponding to the ’t Hooft–Feynman gauge where the NG mode remains explicit. The field $\pi^{\mu}$ is introduced directly into $\beta^{\mu}$ via Eq. (4.2), and therefore represents physical fluctuations of the fixed thermal state.
Evaluating the Lie derivative,
£$\pounds_{\pi}\beta^{\mu}_{c}=\pi^{\rho}\partial_{\rho}\beta^{\mu}_{c}-\beta^{\rho}_{c}\partial_{\rho}\pi^{\mu}=-\beta_{c}\dot{\pi}^{\mu},$ (4.3)
we obtain
$\beta^{\mu}=\beta_{c}(\delta^{\mu}_{0}+\dot{\pi}^{\mu}).$ (4.4)
Under a general coordinate transformation $x^{\mu}\to x^{\mu}-\xi^{\mu}(x)$, we have
$\beta^{\mu}(x)\to\beta_{c}\delta^{\mu}_{0}+\beta_{c}(\dot{\pi}^{\mu}+\dot{\xi}^{\mu})+\mathcal{O}(\xi^{2}).$ (4.5)
Thus, general coordinate transformations are nonlinearly realized on $\pi^{\mu}$:
$\pi^{\prime\mu}(x)=\pi^{\mu}(x)+\xi^{\mu}(x).$ (4.6)
The local temperature, velocity, and chemical potential are defined as
$\beta$ $\equiv\sqrt{-g_{\mu\nu}\beta^{\mu}\beta^{\nu}},$ (4.7)
$u^{\mu}$ $\equiv\frac{\beta^{\mu}}{\beta},$ (4.8)
$\mu$ $\equiv\frac{A_{\mu}\beta^{\mu}}{\beta}.$ (4.9)
For simplicity, we take the background fields as
$g_{\mu\nu}=\eta_{\mu\nu},~A_{\mu}=\mu_{c}\delta^{0}_{\mu}.$ (4.10)
In this case, we find
$\beta$ $=\beta_{c}(1+\dot{\pi}^{0}),$ (4.11)
$u^{\mu}$ $=(1,\dot{\pi}^{i}),$ (4.12)
$\mu$ $=\mu_{c}+\dot{\theta}.$ (4.13)
Thus, hydrodynamical perturbation variables are expressed by the NG modes, and now the covariance becomes manifest. While $\mu$, $\beta$, and $u^{\mu}$ are zeroth-order quantities in the constitutive relations, they are first order in derivatives of the NG modes.
The induced metric $\gamma^{\mu\nu}$, defined with respect to $u^{\mu}$, is distinct from the spatial metric on constant time hypersurfaces. We find
$\gamma^{00}=0,~\gamma^{0i}=\dot{\pi}^{i},~\gamma^{ij}=\delta^{ij}.$ (4.14)
Hydrodynamical coefficients are perturbed via $\beta$ and $\mu$. For example, the energy density $\rho$ expands as
$\rho(\beta,\mu)=\rho_{\rm c}+\beta_{\rm c}\rho_{\beta}\dot{\pi}^{0}+\rho_{\mu}\dot{\theta},$ (4.15)
where $X_{\mu}\equiv(\partial X/\partial\mu)_{\beta}$ and $X_{\beta}\equiv(\partial X/\partial\beta)_{\mu}$ for $X=P,\rho,n$. Hereafter, for notational simplicity, we omit the subscript “c” and redefine the background values as $\beta\equiv\beta_{\rm c}$ and $\mu\equiv\mu_{\rm c}$.
So far, we have only discussed the SSB of the retarded sector in the Schwinger-Keldysh formalism. This sector has advanced counterparts, often referred to as “noise” fields. The Ward identities for the retarded fields,
$\nabla_{\mu}J^{\mu}$ =0=0,= 0 , (4.16)
$\nabla_{\mu}T^{\mu\nu}+J_{\mu}F^{\mu\nu}$ =0=0,= 0 , (4.17)
imply a symmetry of the generating functional under the following transformations:
$\delta_{\bar{\xi},\bar{\Lambda}}g_{\mu\nu}$ =0=0,= 0 , (4.18)
$\delta_{\bar{\xi},\bar{\Lambda}}\bar{g}_{\mu\nu}$ $=\nabla_{\mu}\bar{\xi}_{\nu}+\nabla_{\nu}\bar{\xi}_{\mu},$ (4.19)
$\delta_{\bar{\xi},\bar{\Lambda}}A_{\mu}$ =0=0,= 0 , (4.20)
$\delta_{\bar{\xi},\bar{\Lambda}}\bar{A}_{\mu}$ $=-\nabla_{\mu}\bar{\Lambda}+\bar{\xi}^{\nu}\nabla_{\nu}A_{\mu}+A_{\nu}\nabla_{\mu}\bar{\xi}^{\nu}.$ (4.21)
These transformations reflect redundancies in the doubled degrees of freedom on the CTP. In open systems, this symmetry is generally broken and the energy momentum conservation is not ensured anymore.
One can restore the above symmetries by introducing Stueckelberg fields:
$\bar{G}_{\mu\nu}$ $=\bar{g}_{\mu\nu}-\nabla_{\mu}{\bar{\pi}}_{\nu}-\nabla_{\nu}{\bar{\pi}}_{\mu},$ (4.22)
$\bar{C}_{\mu}$ $=\bar{A}_{\mu}+\nabla_{\mu}\bar{\theta}-{\bar{\pi}}^{\nu}\nabla_{\nu}A_{\mu}-A_{\nu}\nabla_{\mu}{\bar{\pi}}^{\nu}.$ (4.23)
Since SSB occurs in the retarded sector, these Stueckelberg fields can be interpreted as the advanced-sector counterparts of the NG modes.
A more refined analysis is possible by carefully treating the duplicated fields defined on the CTP. To summarize, the set of variables $\Phi=\{\beta,u^{\mu},\mu,F_{\mu\nu},g_{\mu\nu}\}$ and $\bar{\Phi}=\{\bar{G}_{\mu\nu},\bar{C}_{\mu}\}$, along with their covariant derivatives, constitute the symmetry-preserving building blocks for the dynamical action of NG modes.
4.2 Effective action for Nambu-Goldstone modes
@@
When NG modes are set to zero, the effective action must reproduce the UV generating functional (2.8). Hence, the effective action for the NG modes should be written as [24]
$I[\Phi,\bar{\Phi}]=\int d^{4}x\sqrt{-\det g}\sum_{n=1}\mathcal{L}_{n},$ (4.24)
@@
$\mathcal{L}_{1}$ $=\frac{1}{2}T^{\mu\nu}[\Phi]\bar{G}_{\mu\nu}+J^{\mu}[\Phi]\bar{C}_{\mu},$ (4.25)
$\mathcal{L}_{2}$ $=\frac{i}{4}W^{\mu\nu\rho\sigma}[\Phi]\bar{G}_{\mu\nu}\bar{G}_{\rho\sigma}+iX^{\mu\nu\rho}[\Phi]\bar{G}_{\mu\nu}\bar{C}_{\rho}+iY^{\mu\rho}[\Phi]\bar{C}_{\mu}\bar{C}_{\rho}.$ (4.26)
@@
$I_{s}^{(1,0,0)}=\frac{1}{2}\int d^{4}x\Bigg{[}\rho\bar{g}_{00}+P\delta^{ij}\bar{g}_{ij}+2n\bar{A}_{0}\Bigg{]}.$ (4.28)
@@
$\bar{A}_{0}(n_{\mu}\dot{\theta}+\beta n_{\beta}\dot{\pi}^{0})+\bar{A}_{i}(n\dot{\pi}^{i})\Bigg{]}.$ (4.29)
@@
$+\bar{A}_{i}\beta\left(-\beta\chi_{1}\left(\partial^{i}\dot{\theta}+\mu\partial^{i}\dot{\pi}^{0}\right)+\tau_{1}\partial^{i}\dot{\pi}^{0}-\tau_{2}\ddot{\pi}^{i}\right)\Bigg{]}.$ (4.30)
By differentiating Eq. (4.30), one finds the first order fluctuations to $T$ and $J$. The noise term is already presented in Eq. (2.67).
@@
$I_{d}^{(1,0,0)}=-\int d^{4}x\Bigg{[}(\rho-\mu n)\dot{{\bar{\pi}}}_{0}+P\partial^{i}{\bar{\pi}}_{i}-n\dot{\bar{\theta}}\Bigg{]}.$ (4.31)
@@
$I^{(1,0,0)}_{d}=\int d^{4}x\Bigg{[}(\dot{\rho}-\mu\dot{n}){\bar{\pi}}_{0}+\dot{n}\bar{\theta}\Bigg{]}.$ (4.32)
@@
$+n\partial_{i}\bar{\theta}\dot{\pi}^{i}+\dot{\bar{\theta}}\left(\dot{\theta}n_{\mu}+\beta\dot{\pi}^{0}n_{\beta}\right)\Bigg{]}.$ (4.34)
@@
$+\partial_{i}\bar{\theta}\beta\left(-\beta\chi_{1}\left(\partial^{i}\dot{\theta}+\mu\partial^{i}\dot{\pi}^{0}\right)+\tau_{1}\partial^{i}\dot{\pi}^{0}-\tau_{2}\ddot{\pi}^{i}\right).$ (4.35)
@@
$-2\nu_{1}\dot{\bar{\pi}}_{0}(\dot{\bar{\theta}}+\mu\dot{\bar{\pi}}_{0})-2\tau\left(\partial_{i}{\bar{\pi}}_{0}+\dot{{\bar{\pi}}}_{i}\right)(\partial^{i}\bar{\theta}+\mu\partial^{i}\bar{\pi}_{0})-2\nu_{2}(\dot{\bar{\theta}}+\mu\dot{\bar{\pi}}_{0})\partial^{i}\bar{\pi}_{i}\Bigg{]}.$ (4.36)
@@
Before going to the mode analysis, let us explain the present expansion scheme. In the first-order hydrodynamical theory, there are two energy scales: $M_{\rm SSB}$ and $M_{\rm diss.}$.
$M_{\rm SSB}$ is the energy of spontaneous symmetry breaking, typically $\rho+P\sim M_{\rm SSB}^{4}$, which is the decay constant in nonlinear realization theory. For a conformal fluid, $M_{\rm SSB}\sim\beta^{-1}$. The perturbative description is not applicable for $\omega>M_{\rm SSB}$, which is typically the energy scale of a thermal fast mode. In fact, one can canonically normalize the NG mode as $\pi_{\rm c}\sim\sqrt{\rho+P}\pi$, and the action is expanded by $(\partial\pi_{\rm c}/M^{2}_{\rm SSB})^{n}$. Hence, the perturbative expansion is justified for $\omega\ll M_{\rm SSB}$. $\eta\sim M_{\rm diss.}^{3}$ is the energy scale of dissipation determined by the mean-free-path (MFP) of the fast modes. The order counting parameter of the gradient expansion is $\varepsilon\sim\omega M_{\rm diss.}^{3}/M_{\rm SSB}^{4}$, and the gradient expansion is valid in the regime where $\varepsilon\ll 1$.
Section summary.
In this section, we argued that the spontaneous breaking of spacetime symmetry occurs in the presence of a local thermal state characterized by the local four-vector $\beta^{\mu}$. We then demonstrated that perturbations in $\beta^{\mu}$ can be expressed in terms of NG modes, analogous to the embedding of NG modes in global U(1) theories. The spacetime NG mode $\pi^{\mu}$ nonlinearly realizes diffeomorphism symmetry, enabling a manifestly covariant formulation of hydrodynamical perturbations. We also introduced the noise counterparts of the NG modes and constructed the corresponding hydrodynamical effective action. The remaining task is a straightforward, albeit technically involved, mode analyis of this effective action.